\begin{thebibliography}{99}
\bibitem[CaSc22]{CaSc22} S.~L.~Cacciatori and A.~Scotti,
``Compact Lie Groups, Generalised Euler Angles, and Applications,''
Universe \textbf{8} (2022) no.10, 492
\bibitem[CacUrs22]{CacUrs22} S.~L.~Cacciatori and P. Ursino, ``Concentration of measure for classical Lie groups" European Journal of Mathematics {\bf 9}, no. 3 Springer(2023)
https://doi.org/10.1007/s40879-023-00607-2.
\bibitem[CU22]{CU22}
 S.~L.~Cacciatori, P.~Ursino,
``Macdonald formula, Ricci Curvature, and Concentration Locus for classical compact Lie groups,''  Axioms 2022, 11(6), 245; https://doi.org/10.3390/axioms11060245 .
\bibitem[CanSin]{CanSin} ~P.~Cannarsa
~C.~Sinestrari, ``Semiconcave Functions,
Hamilton–Jacobi Equations,
and Optimal Control," Birkhauser, Berlin, (2004).
\bibitem[CDPS]{CaDaPiSc}S.~L.~Cacciatori, F.~Dalla Piazza, A.~Scotti,
``Compact Lie groups: Euler constructions and generalized Dyson conjecture,''
Trans. Am. Math. Soc. {\bf 369}, no. 7 (2017), 4709--4724.
\bibitem[CheeEbin]{CheeEbin1} J. Cheeger and D. Ebin, ``Comparison Theorems in Riemann Geometry," Mathematical Library, North Holland Elsevier, Amsterdam, (1975).

\bibitem[Fath]{Fath}  ~A.~ Fathi
https://www.math.u-bordeaux.fr/~pthieull/Recherche/KamFaible/Publications/Fathi200801.pdf

\bibitem[FigRiffVil]{FigRiffVil}
Figalli, A., Rifford, L. and Villani, C.
"Nearly round spheres look convex"
Amer. J. Math. {\bf 134} (2012), no. 1, 109–139.
\bibitem[G]{Gl} H. Gl\"ockner, ``Fundamentals of direct limit Lie theory," Compositio Mathematica, {\bf 141} no. 6 (Nov., 2005),
1551--1577.
\bibitem[GM]{GrMi}
 M.~Gromov and V.~D.Milman,
 ``A Topological Application of the Isoperimetric Inequality,''
 American Journal of Mathematics, {\bf 105} no. 4 (Aug., 1983), 843--854.
\bibitem[LD]{ledoux} M. Ledoux. ``The Concentration of Measure Phenomenon,''
American Mathematical Society (2001), USA

\bibitem[Gray]{Gray}
 Alfred Gray, "Tubes", Progress in Mathematics
221, Birkhauser Boston, Inc., Basel, MA, 2004.


\bibitem[Gro99]{Gro99}
 Michail Gromov, "Metric structures for Riemannian and non-Riemannian spaces", Progress in Mathematics
152, Birkhauser Boston, Inc., Boston, MA, 1999.



\bibitem[ItoTanaka]{ItoTanaka} Jin-Ichi Itoh, Minoru Tanaka, ``The Dimension Of A Cut Locus On A Smooth
Riemannian Manifold" Tohoku Math. J. {\bf 50} (1998), 571--575.

\bibitem[LOE]{LOE} Loeper, Grégoire "Regularity of optimal maps on the sphere: the quadratic cost and the reflector antenna" Arch. Ration. Mech. Anal. {\bf 199} (2011), no. 1, 269–289.

\bibitem[LV]{LoVi} J.Lott and C.Villani, ``Ricci curvature for metric-measure spaces via optimal transport," Annals of Mathematics, {\bf 169} (2009), 903--991.

\bibitem[ManteMen]{ManteMen}
~C.~Mantegazza and ~A.~C.~G.~Mennucci, "Hamilton-Jacobi Equations and Distance Functions on Riemannian Manifold" Applied Mathematics and Optimization.
{\bf 47}. no.1 (2003), 1--25.


\bibitem[McCannFeld]{MCannFeld} Mikhail Feldman And Robert J. Mccann, ``Monge's transport problem
on a riemannian manifold" Transactions of the AMS {\bf 354} no.4 (2001) 1667–1697.

\bibitem[McCannPass]{McCannPass} Brendan Pass and Robert J. Mccann, ``Optimal transportation between
unequal dimensions" Arch. Ration. Mech. Anal. {\bf 238} (2020), no. 3, 1475–1520.
\bibitem[McCannYoung]{McCannYoung}
Kim, Young-Heon and McCann, Robert J. "Towards the smoothness of optimal maps on Riemannian submersions and Riemannian products (of round spheres in particular)" J. Reine Angew. Math. {\bf 664} (2012), 1–27.
\bibitem[M]{Ma}
I. G. Macdonald, ``The Volume of a Topological Group,'' Invent. Math. {\bf 56} no. 93 (1980).

\bibitem[P1]{Pe1} V. Pestov, ``Dynamics of Infinite-Dimensional Groups: The Ramsey-Dvoretzky-Milman Phenomenon," University Lecture
Series {\bf 40}, American Mathematical Society, Providence, RI, (2006).
\bibitem[S]{Sc} F. M.Schneider, ``Equivariant Concentration In Topological Groups," Geometry \& Topology {\bf 23} no.2 (2019) 925–956.


\bibitem[Shi]{Shi} Takashi Shioya, Metric measure geometry: Gromov's theory of convergence and concentration of
metrics and measures, IRMA Lectures in Mathematics and Theoretical Physics 25, EMS Publishing House, Zurich, 2016.

\bibitem[ShiKa]{ShiKa}Daisuke Kazukawa, Takashi Shioya, ``High-Dimensional Ellipsoids
converge to Gaussian Spaces," Arkiv 2003.05.105v [MATH MG].

\bibitem[VilBook1]{VilBook1} Topics in optimal transportation, Graduate Studies in Mathematics, vol. 58, American
Mathematical Society, Providence, RI, 2003

\bibitem[VilBook]{VilBook} C.Villani ``Optimal Transport, old and new''
Springer and Verlag (2009), Berlin.

\bibitem[We]{Weyl} H. Weyl


\end{thebibliography}

\documentclass[11pt,reqno]{amsart}

\usepackage{amsmath,amssymb}
\usepackage{latexsym}
\usepackage{epsfig}
\usepackage{cite}
\usepackage[english]{babel}
\usepackage{booktabs}
\usepackage{graphicx}
\usepackage{tikz}
\usepackage{pgfplots}
\usepackage{eurosym}
\usepackage[inner=4cm,outer=2cm]{geometry}
\usepackage{amsthm}
\usepackage{physics}

\usetikzlibrary{arrows,decorations.pathmorphing,backgrounds,positioning,fit,petri,decorations}
\usetikzlibrary{calc,intersections,through,backgrounds,mindmap,patterns,fadings}
\usetikzlibrary{decorations.text}
\usetikzlibrary{decorations.fractals}
\usetikzlibrary{fadings}
\usepackage{pgflibraryshapes}
\usetikzlibrary{lindenmayersystems}
\usetikzlibrary{shadings}
\usetikzlibrary{shadows}
\usetikzlibrary{shapes.geometric}
\usetikzlibrary{shapes.callouts}
\usetikzlibrary{shapes.misc}
\usetikzlibrary{spy}
\usetikzlibrary{topaths}
\usetikzlibrary{datavisualization}
\usetikzlibrary{datavisualization.formats.functions}

\newenvironment{sistema}
{\left\lbrace\begin{array}{@{}l@{}}}
{\end{array}\right.}

\newtheorem{teo}{Theorem}
\newtheorem{rem}{Remark}
\newtheorem{prop}{Proposition}
\newtheorem{cor}{Corollary}
\newtheorem{tdef}{Definition}

\numberwithin{equation}{section}

\numberwithin{equation}{section}

\title{Volume of Tubes and Concentration of Measure in Riemannian Geometry}
\author{S.L.~Cacciatori$^{1,2}$ and P.~Ursino$^{3}$}

\address{$^1$ Department of Science and High Technology, Universit\`a dell'Insubria, Via Valleggio 11, IT-22100 Como, Italy}
\address{$^2$ INFN sezione di Milano, via Celoria 16, IT-20133 Milano, Italy}
\address{$^3$ Department of Mathematics and Informatics, Universit\`a degli Studi di Catania, Viale Andrea Doria 6, 95125 Catania, Italy}

\begin{document}
\maketitle
\begin{abstract}

We investigate the notion of concentration locus introduced in \cite{CacUrs22}, in the case of Riemann manifolds sequences and its relationship with the volume of tubes. After providing a general formula for the volume of a tube around a Riemannian submanifold of a Riemannian manifold, we specialize it to the case of totally geodesic submanifolds of compact symmetric spaces. In the case of codimension one, we prove explicitly concentration. Then, we investigate for possible characterizations of concentration loci in terms of Wasserstein and Box distances.

\end{abstract}
\section*{Introduction}
A concentration locus is roughly speaking a sequence of sub-manifolds $(M_n,\sigma_n,g_n)$ (where $g_n$ is the geodetic metric and $\mu_n$ the volume measure) which approximates the concentration behaviour of the manifolds $(N_n,\mu_n,g_n)$ where they are embedded.
In a sense, the concentration character of the ``big'' sequence is fully determined by the ``thin'' one.
This phenomenon is particularly significant, from the point of view of applications, whenever it is possible to single out, inside a sequence of manifolds of unknown concentration behaviour, a sequence of much simpler sub-manifolds which is a concentration locus and therefore, provided the concentration behaviour of the sub-manifolds is known, it determines the concentration behaviour of the big one.

Both sequences can be regarded as sequences of metric-measurable spaces (mm-spaces). In the space $\mathfrak{M}$ of all mm-spaces Gromov, in his celebrated green book \cite{Gro99}, defines the notion of observable distance ($d_{conc}$ in \cite{Shi}) which fully generalizes the classical phenomenon of concentration of measure to a point and Levy family (see for example \cite{GrMi}).
Practically, we say that a sequence of mm-spaces concentrates to an mm-space, whenever the former $d_{conc}$ converges to the latter.

Our primary goal consists in finding a way to detect if a sequence of sub-manifolds is or is not a concentration locus by investigating the sequences of volumes of tubes built around the sub-manifolds, with decreasing rays.

The problem of calculating the volume of a tube in a Riemann manifold is very interesting in itself and it is treated in Section 1. We start from the well-known article of Weyl \cite{Weyl} and the beautiful book of Gray \cite{Gray}.
We succeed, using the approach of Gray, in finding a general formula for the volume of tubes which involves the codimension, the ray of the tube, the curvature of the ambient manifold and the killing curvatures of the sub-manifold.
Unfortunately, it is widely useless in computing concentration locus unless you have an estimate of the curvature derivatives of every order.
The case of symmetric spaces is much more practicable, instead. Indeed, we find a concrete formula and we are able to extend the system of coordinates used for calculating the volume of the tubes to the entire ambient manifold and therefore calculate the asymptotic conditions to detect a concentration locus.  We believe that our results can be easily extended to the case of homogeneous spaces.

Observable distance is a very difficult tool to deal with, fortunately, there are more practical distances that are related to $d_{conc}$.
For example, Wasserstein distance $d_{W}$, which is related to optimal transport, and $d_{box}$, which makes $\mathfrak{M}$ a complete metric space. The following holds $d_{W}\Rightarrow d_{box}\Rightarrow d_{conc}$ even if $d_{W}\nLeftarrow d_{box}\nLeftarrow d_{conc}$ (there are $d_{conc}$ cauchy sequence which does not converge \cite{ShiKa}).
On the other side, it is very hard to find differentiable optimal transport. Indeed,  McCann et al. showed that the Monge-Kantorovich problem is solvable for smooth Riemannian manifolds \cite{MCannFeld}. Regarding differentiable solutions, Figalli, Rifford, and Villani \cite{FigRiffVil} solved it in the positive for the spheres \cite{FigRiffVil}, while McCann and Young-Heon for $\mathbb {CP}^n$ \cite{McCannYoung}.

However, our Wasserstein problem is a special one, since we are looking for a transport to the push forward measure induced by the transport itself. In this case, the problem assumes differential geometry features and a viscosity solution of the following Hamilton-Jacobi equation
\[
 \begin{sistema}
	|\nabla u |^2=1\ \mbox{in} \ M\setminus C\\
	u=0\ \mbox{in} \ C
\end{sistema}
\]
provides an answer to our problem. In particular, the distance function, which solves the previous equation, singles out a projection function which is exactly the function that minimizes the distance from the target sub-manifold. This solution is unique up to a measure null set (focal points) and up to this set it is differentiable.

By a well-known result, which is a consequence of Hopf-Rinow Theorem (see Proposition 5 Cheeger \cite{CheeEbin1}), the trajectories (the geodesics) of this projection, through which the mass is transported, are orthogonal to the boarder of $C$. In particular,
$\gamma$, the geodesic or trajectory which has length exactly the distance, is uniquely determined and,
it is the super-differential of the distance function.

All these results can be seen in \cite{CanSin} in $R^n$ context or in \cite{ManteMen} or in \cite{Fath} in the Riemann geometry context.

We succeed in Section 2 in finding a characterization of concentration locus in terms of $d_{box}$ and projection distance, and this allows us to prove Corollary \ref{ConcLocConv}, which guarantees that whenever a concentration locus converges $d_{conc}$ to an mm-space $(H,\mu,g)$, the sequence of ambient manifolds concentrates to $(H,\mu,g)$.

\section*{Preliminaries.}
Given a set $X$, we denote by $l^{\infty}(X)$ the unital
Banach algebra of all bounded real-valued functions on $X$ equipped with the supremum norm.
Let $X$ be a topological space. If the topology of $X$ is generated by a metric $d$, then we call
$d$ a compatible metric on $X$.\\
We will denote by $C(X)$ the set of all continuous real-valued functions on $X$, and we set $CB(X)=l^{\infty}(X)\cap C(X)$. Let us denote by $\mathcal{B}(X)$ the
Borel $\sigma$-algebra of $X$ and by $P(X)$ the set of all Borel probability measures on $X$.
The weak topology on $P(X)$ is defined to be the initial topology on $P(X)$ generated by the maps of the
form $P(X) \rightarrow \mathbb{R}, \mu\rightarrow \int f d\mu$ where $f \in CB(X) $.\\
The support of a measure $\mu\in P(X)$ is defined as
\begin{align*}
spt\mu=\{x\in X\mid \forall U\subseteq X\mbox{ open: }x\in U\Rightarrow \mu (U)>0\}.
\end{align*}
Given $\mu\in P(X)$ and a Borel subset $B \subseteq X$
with $\mu(B) = 1$, we let $\mu|_B :=\mu|_{\mathcal{B}(B)\subset P(B)}$. The push-forward of a measure $\mu\in P(X)$ along
a Borel map $f : X\rightarrow Y$ into another topological space $Y$ is defined to be
\begin{align*}
 \sharp f\mu:\mathcal{B}(Y)\rightarrow [0,1], B\mapsto \mu (f^{-1}(B)).
\end{align*}
Furthermore, let us note that each $\mu\in P(X)$ gives rise to a pseudo-metric
$me_{\mu}$ on the set of
all Borel measurable real-valued functions $X$ defined by
\begin{align*}
me_{\mu}(f,g):=\inf\{\epsilon >0 \mid \mu (\{x\in X\mid \mid f(x)-g(x)\mid >\epsilon \})\le\epsilon\},
\end{align*}
for any two Borel functions $f,g:X\rightarrow\mathbb{R}$.
Let $(X, d)$ be a pseudo-metric space. Given a subset $A\subseteq X$, we abbreviate $d|_A := d|_{A\times A}$ and define ${\rm diam}(A,d):=\sup\{d(x,y)\mid x,y\in A\}$. For $x\in A\subseteq X$ and $\epsilon >0$ we set
\begin{align*}
B_d(x,\epsilon ):=\{y\in X\mid d(x,y)<\epsilon \}\ \ B_d(A,\epsilon ):=\{y\in X\mid \exists a\in A ; d(x,y)<\epsilon \}.
\end{align*}
Then the Hausdorff distance between any two subsets $A,B \subseteq X$ is given by
\begin{align*}
d_H(A,B):=\inf \{\epsilon >0\mid B\subseteq B_d(A,\epsilon ),\ A\subseteq B_d(B,\epsilon )\}.
\end{align*}
For $l,r\ge 0$, we denote by $Lip_l(X, d)$ the set of all $l$-Lipschitz real-valued functions on $(X, d)$,
and we define
\begin{align*}
Lip_l^{\infty}(X,d):=Lip_l(X,d)\cap l^{\infty}(X),\ Lip_l^{r}(X,d):=\{f\in Lip_l(X,d)\mid \|f\|_{\infty}\le r\}.
\end{align*}
Moreover, we set $Lip(X,d):=\bigcup \{Lip_l(X,d)\mid l\ge 0\}$ and $Lip(X,d)^{\infty}:=Lip(X,d)\cap l^{\infty}(X)$.\\
Whenever $(X,d)$ is a separable metric space, the Wasserstein distance $W_1(\mu,\nu)$\footnote{Different names appearing in the literature include Monge-Kontorovich distance, bounded Lipschitz distance,
mass transportation distance, and Fortet-Mourier distance \cite{VilBook}} is a compatible metric for a weak topology on $P(X)$ defined by
\begin{align*}
W_1(\mu,\nu):=\sup_{f\in Lip_1^{1}(X,d)}\left | \int f d\mu -\int f d\nu \right | \ \ (\mu.\nu\in P(X)).
\end{align*}

\begin{tdef}{(Gromov-Milman \cite{Gro99})}\\
  A space with a metric $g$ and a measure $\mu$, or an mm-space, is a triple $(X,\mu,g)$, consisting of a set $X$, a metric $g$ on $X$ and a probability Borel measure such that $(X,g)$ is a separable complete metric space.
\end{tdef}
Moreover, an mm-space $(X,\mu ,d)$ is called compact if $(X, d)$ is compact, and fully supported if $spt \mu= X$.
Henceforth, we will denote by $\lambda$ the Lebesgue measure on $[0, 1)$.\\
A parametrization of an mm-space $(X,\mu ,d)$ is a Borel measurable map $\phi: [0, 1)\rightarrow X$ such that $\sharp\phi\lambda = \mu$. It is well known that any mm-space admits a parametrization (see, e.g.\cite{Shi}).
In the set of isomorphism classes of mm-spaces, $\mathfrak{M}$, we can define the box distance, $d_{box}$, that we are going to define (\cite{Shi}).
For two pseudo-metrics $\rho_1$ and $\rho_1$
on the unit interval $I$, we define $d_{box}(\rho_1,\rho_2)$ to be the infimum of $\epsilon >0$ satisfying that
there exists a Borel subset $I_0\subseteq I$ such that
\begin{enumerate}
          \item $\vert\rho_1(s,t),\rho_2(s,t)\vert\le\epsilon$ for any $s,t\in I_0$,
          \item $\mathfrak{L}^1(I_0)\ge 1-\epsilon$ where $\mathfrak{L}^1$ denotes the one-dimensional Lebesgue measure.
\end{enumerate}

\begin{tdef}
  let $X$ be a topological space with a Borel probability measure $\mu_X$. A map $\varphi:I\rightarrow X$ is called a parameter of $X$ if $\varphi$ is a Borel measurable map such that

  $\sharp\varphi\mathfrak{L}^1=\mu_X$
\end{tdef}

We define box distance $d_{box}$ between two isomorphism classes of mm-spaces $X,Y$ to be the infimum of $d_{box}(\sharp\varphi d_X,\sharp\psi d_Y)$ where
$\varphi:I\rightarrow X$ and $\psi:I\rightarrow Y$ run over all parameters of $X$ and $Y$, respectively, and where $\sharp\varphi d_X(s,t):=d_X(\varphi(s),\varphi(t))$.

\begin{tdef}{(Gromov-Milman \cite{Gro99}\cite{Shi})}\\
In the set of isomorphism classes of mm-spaces we can define the following distance:
\begin{align*}
d_{conc}(X,Y):= \inf\{(me_{\lambda})_H(Lip_1(X)\circ\phi , Lip_1(Y)\circ\psi\mid \phi\mbox{ param. of }X,\psi\mbox{ param. of }Y\}.
\end{align*}
\end{tdef}
Two mm-spaces $X$ and $Y$ are isomorphic if there exists an isomorphism between mm-spaces $(X,\mu ,d),(Y,\nu ,d')$ i.e an isometry
\begin{align*}
f: (spt \mu,d\mid X)\rightarrow (spt \nu,d'\mid Y)
\end{align*}
such that  $\sharp f(\mu\mid {spt\mu})=\nu\mid {spt\nu}$.
A sequence of mm-spaces $(X_n,\mu_n,g_n)$ is said to concentrate to an mm-space $(X,\mu,g)$ if
\begin{align*}
\lim_n d_{conc}(X_n,X)=0.
\end{align*}

In this case, we denote $(X,\mu,g)$ as a concentration set for the sequence of mm-spaces $(X_n,\mu_n,g_n)$.\\
Finally, let us recall the definition of Concentration Locus as defined in \cite{CacUrs22}

\begin{tdef}\label{def2}
 Let $\{X_n, \mu_n\}_{n\in \mathbb N}$ be a family of metric spaces with metrics $d_n$, and Borel's measures $\mu_n$ w.r.t. which nonempty open set have non-vanishing measure. Assume the measures to be normalized, $\mu_n(X_n)=1$.
 Let $\{S_n\}_{n\in \mathbb N}$ be a family of proper closed subsets, $S_n\subset X_n$. Fix a sequence $\{\varepsilon_n\}_{n\in \mathbb N}$ such that $\varepsilon_n>0$,  $\lim_{n\to\infty}\varepsilon_n=0$, and let
 $\{U^{\varepsilon_n}_n\}_{n\in \mathbb N}$ be the sequence of tubular neighbourhoods of $S_n$ of radius $\varepsilon_n$. We say that the family
 $\{S_n\}$ is a Concentration Locus if
 \begin{align}
 \lim_{n\to\infty} \mu_n(X_n-U^{\varepsilon_n}_n)=0.
 \end{align}

 Moreover, if such a sequence $\varepsilon_n$ converges to 0 at rate $k$ (so that $\lim_{n\to\infty} n^k\varepsilon_n=c$ for some constant $c$), we say that the family $\{S_n\}$ is a Concentration Locus at least at rate $k$.
\end{tdef}

\section{Volume of tubes in compact manifolds. }
Let $M$ be a compact Riemannian submanifold $M\subseteq N$ of codimension $q$ of a manifold $N$. We call $M_{\epsilon}$ the tube generated by all geodetic segments of length $\epsilon$, outgoing perpendicularly from $M$. It is a well-known result established by Weyl that the volume of the tube can be expressed in terms of the Lipschitz-Killing curvatures $K_{2j}$ of $M$ and the ambient space, and the codimension $q$. More precisely:
\begin{teo}[Weyl, \cite{Weyl}]
 Let $M$ be a compact Riemannian submanifold of $\mathbb R^N$ of codimension $q=N-n$. Let $M_\varepsilon$ a tubular neighbourhood of $M$ of radius $\varepsilon$. Then, for all $r>0$ sufficiently small, it holds
 \begin{align}
 {\rm Vol}_{\mathbb R^N}(M_\varepsilon)=\frac {\pi^{\frac q2}\varepsilon^q}{\Gamma(\frac q2+1)} \left(K_0(M) +\sum_{j=1}^{\lfloor n/2\rfloor} \frac {K_{2j}(M)\varepsilon^{2j}}{(q+2)(q+4)\cdots (q+2j)}\right),
 \end{align}
 where
 \begin{align}
 K_{2j}(M) =\int_M k_{2j}(\Omega),
 \end{align}
 are the integrated Lipschitz-Killing curvatures, and $\Omega$ is the curvature 2-form of $M$.
\end{teo}
Remember that if $e^a$ is the dual basis to an orthonormal frame $V_a$, $a=1,\ldots,n$, then
\begin{align}
 k_{2j}(\Omega)=\frac 1{2^j j! (n-2j)!} \sum_{\sigma\in S_n} \epsilon_\sigma \Omega_{\sigma(1)\sigma(2)}\wedge \cdots \wedge \Omega_{\sigma(2j-1)\sigma(2j)} e^{\sigma(2j+1)}\wedge \cdots \wedge e^{\sigma(n)},
\end{align}
where $S_n$ is the set of permutations of $n$ elements and $\epsilon_\sigma$ the sign of the permutation. In particular,
\begin{align}
 k_0(\Omega)=d{\rm Vol}_M,
\end{align}
is the volume form on $M$,
\begin{align}
 k_2(\Omega)=&\frac 12 R\ d{\rm Vol}_M, \\
 k_n(\Omega)=& {\rm Pf}(\Omega),
\end{align}
where $R$ is the scalar curvature of $M$ and Pf the Pfaffian. Finally, notice that
\begin{align}
 \frac {\pi^{\frac q2}\varepsilon^q}{\Gamma(\frac q2+1)}={\rm Vol}_{\mathbb R^q} (D_\varepsilon)
\end{align}
is the volume of the $q$ dimensional disc of radius $\varepsilon$ in $\mathbb R^q$. So, if we define the {\it mean Lipschitz-Killing curvatures} $\kappa_{2j}$ as
\begin{align}
\kappa_{2j}=\frac {K_{2j}(M) }{{\rm Vol}_M(M)},
\end{align}
then we can rewrite Weyl's formula as
 \begin{align}
 {\rm Vol}_{\mathbb R^N}(M_\varepsilon)={\rm Vol}_M(M) {\rm Vol}_{\mathbb R^q} (D_\varepsilon)  \left(1 +\sum_{j=1}^{\lfloor n/2\rfloor} \frac {\kappa_{2j}(M)\varepsilon^{2j}}{(q+2)(q+4)\cdots (q+2j)}\right).\label{formula}
 \end{align}
We are interested in understanding the volumes of tubular neighborhoods of submanifolds of compact manifolds (and, in general, on manifolds with positive curvature). However, it is interesting to do some general considerations on this formula before discussing
the more general case.

\

Since there is no curvature in the directions of $\mathbb R^N$ orthogonal to $M$, the deformations of the volume are only due to the bending of $M$. If $M$ is flat, then the volume of the tube is just the product of the volumes of the submanifold and the disc.
The terms in the parenthesis then give the contributions of the deformations of the tube to the volume, when we bend the tube along a curved $M$. For example, if we bend the tube neighborhood of a segment to the one of a circle, the tube will be compressed
along the most internal circle and stretched along the most external one. However, the volume doesn't change (if we don't change the length of the segment), indeed the scalar curvature of the circle is $R_{S^1}=0$.\\
Now, since $\kappa_{2j}$ are mean curvatures, we can get some hints about the dependence on curvatures by considering $M$ to be a manifold of constant sectional curvature $1/r^2$. In this case, one has
\begin{align}
 \varepsilon^{2j} \kappa_{2j}(\Omega)=\frac {n!}{2^j j! (n-2j)!} \left(\frac {\varepsilon}{r}\right)^{2j}.
\end{align}
We want to see under which conditions the curvature terms become relevant in the formula (\ref{formula}) when $n$ grows. In general, this also could imply that generically also $q$ grows. Keeping $j$ fixed, we see that Stirling's formula implies
\begin{align}
 \varepsilon^{2j} \kappa_{2j}(\Omega)\approx \frac 1{2^jj!} \left(\frac {n\varepsilon}{r}\right)^{2j}.
\end{align}
If $\varepsilon/r$ is small but constant, the curvature terms become dominant for large $n$, at least if $q$ is constant. Despite we are considering spheres embedded in $\mathbb R^{n+q}$, let us for a moment imagine assuming $q=1$
and that $\mathbb R^{n+q}$ is replaced by $S_r^{n+q}$. In this case, it is well known that the measure of the whole sphere concentrates in a tube of radius $\varepsilon\sim r n^{-\frac 12}$ around the equator $M$. In this situation, we see that
\begin{align}
 \varepsilon^{2j} \kappa_{2j}(\Omega)\approx \frac 1{2^jj!} n^j,
\end{align}
so the curvature terms are dominant w.r.t. the 1. If also the codimension $q\equiv q_n$ increases unboundedly, then, including the denominators, we see that the contributions are of the order
\begin{align}
 \frac 1{2^jj!} (n/q_n)^j,
\end{align}
so that the dominance of the curvatures persists if $q_n$ grows slower than $n^{1-a}$ for any fixed arbitrarily small but positive $a$. This is also another well-known condition for concentration. Why should we consider these considerations acceptable if we
replace the flat ambient space with spheres? The reason is that the spheres have curvature $1/r^2$ much smaller than the inverse square radius of the tube, $\sim n/r^2$. This suggests that in general, we may obtain the same results if we have an estimation of the bound of the curvature of the ambient manifold $N$ and a control on the error we make in using the flat formulas as a function of the radius of the tube.

\

Since we are interested in compact manifolds, we recall that Weyl deduced the exact formula for the case when the embedding space is a sphere. From this result we can deduce:
\begin{prop}
 Let $M\subset S_R^{n+q}$ a $n$ dimensional smooth compact submanifold of codimension $q$ of a sphere of radius $R$, and $M_\varepsilon$ the tube of radius $\varepsilon$ around $M$. Then, we can write
 \begin{align}
 {\rm Vol}_{S_R^{n+q}}(M_\varepsilon)={\rm Vol}^{flat}(M_\varepsilon) \left(1+o(\varepsilon/R) \right),
 \end{align}
 where ${\rm Vol}^{flat}_N(M_\varepsilon)\equiv {\rm Vol}_{\mathbb R^{n+q}}(M_\varepsilon)$ is given by Weyl's formula for the embedding in the flat $\mathbb R^{n+q}$.
\end{prop}
{\bf Remark:} {\it Before giving the proof, let us notice that this proposition has an immediate consequence: if $\varepsilon\ll R$ then we can use the result of the previous section to deduce the properties of concentration of the measure around $M$, up to a relative error controlled
by $\varepsilon/R$.}
\begin{proof}
 Let us set $N=n+q$. In \cite{Weyl}, Weyl deduced the following formula for the volume of a tube of ``radius $a$'' in a sphere of radius $R$:
 \begin{align}
 {\rm Vol}_{S^N}(M_\varepsilon)=\frac {2\pi^{\frac q2}}{\Gamma(\frac q2)}\sum_{j=0}^{\lfloor n/2\rfloor} \frac {K_{2j}(M) R^{2j+q}}{q(q+2)\cdots (q+2j-2)} \int_0^{\frac \varepsilon{R}} (\sin\rho)^{q+2j-1}(\cos\rho)^{n-2j} d\rho.
 \end{align}
 Here, the radius $a$ is not the geodesic radius along the sphere but the Euclidean radius in the tangent space. It is related to the geodesic radius $\varepsilon$ through the relation $a=R\tan \frac {\varepsilon}R$.
 If we use the change of variable $x=\sin^2\rho/\sin^2(\varepsilon/R)$ we get
 \begin{align}
 \int_0^{\frac \varepsilon{R}} (\sin\rho)^{q+2j-1}(\cos\rho)^{n-2j} d\rho=& \frac 12 (\sin(\varepsilon/R))^{q+2j} \int_0^1 dx\ \frac {x^{\frac q2+j-1}}{\left(1-\sin^2 (\varepsilon/R)x\right)^{j-\frac {n-1}2}}\cr
 =&\frac {(\sin(\varepsilon/R))^{q+2j}}{q+2j} {}_2F_1\left(j+\frac q2,j-\frac {n-1}2;j+\frac q2; \sin^2 \frac {\varepsilon}R\right).
 \end{align}
Thus,
 \begin{align}
 {\rm Vol}_{S^N}(M_\varepsilon)=\frac {2\pi^{\frac q2}}{\Gamma(\frac q2)}\sum_{j=0}^{\lfloor n/2\rfloor} \frac {K_{2j}(M) R^{2j+q}(\sin(\varepsilon/R))^{q+2j}}{q(q+2)\cdots (q+2j)}
 {}_2F_1\left(j+\frac q2,j-\frac {n-1}2;j+\frac q2; \sin^2 \frac {\varepsilon}R\right).
 \end{align}
Using that $\sin x =x+o(x)$ and that ${}_2F_1(a,b;c;x^2)=1+o(x)$ for $x\to 0$, and that the volume of the discs on a sphere are the same as in the tangent space up to correction of order $\varepsilon/R$, we get the proof of our assert.
\end{proof}
This corroborates our previous discussion, showing that we can use formula \eqref{formula} when the radius of the tube is small compared with the radius of curvature of the sphere. Notice that the factor containing the mean curvatures seems to suggest that
one should look for submanifolds having large curvatures in order to look for concentration phenomena. However, it is well-known that in the case of spheres, the concentration is on equators, which are totally geodesic submanifolds with the
property that the extrinsic curvature vanishes. Therefore, in this case, the Lipschitz-Killing curvatures are all zero, contradicting our intuition. Before explaining why this happens, let us see how things work in the more general case.

\subsection{The general case}
Let $M$ be a Riemannian submanifold of codimension $q$ in a compact Riemannian manifold $N$ of dimension $q+n$. Let $M_\varepsilon$ a tube of radius $\varepsilon$ around $M$. We can coordinatize the tube as follows. Fix a point $p$ on $M$ with local coordinates $\bar x$. Consider
the normal bundle $\mathcal NM$ of $M$ in $\Sigma$ and let $\hat n_j(\bar x)$, $j=1,\ldots, q$, an orthonormal basis of $\mathcal N_pM$. We can introduce polar coordinates $\theta_a$, $a=1,\ldots,q-1$,
and director cosines $\omega^j(\bar\theta)$ to parametrise the
arbitrary direction orthogonal to $M$ as $n(\bar \theta;\bar x)=\sum_j \omega^j(\bar\theta)\hat n_j(\bar x)$. Let us consider the geodesic $\gamma_{\bar x, \bar \theta}(t)$ in $N$ starting at $t=0$ from $p$ in the direction $n(\bar \theta;\bar x)$, where $t$ is the geodesic length parameter. The tube of
radius $\varepsilon$ is defined by all such geodesics for $t\leq \varepsilon$. For $\varepsilon$ small enough it is well defined. We use the coordinates $(\bar x, \bar \theta, t)$ in the tube. In order to compute the volume of the tube in the
measure of $N$, we need to compute the measure in the given local coordinates. We do it assuming that the coordinate $\underline x$ cover the whole $M$ up to a subset of vanishing measure. This is not a restriction since this hypothesis can be replaced
by the introduction of a partition of unity.
In $p\equiv \bar x$, let us choose an orthonormal frame in $T_pM$, say $e_a$, $a=1,\ldots,n$, to be used as a frame of Fermi along the normal geodesic $\gamma_{\bar x, \bar \theta}(t)$. The Jacobian we are interested in is
\begin{align}
 J=(\partial_{\bar x} \gamma_{\bar x, \bar \theta}; \partial_{\bar\theta} \gamma_{\bar x, \bar \theta}; \partial_t \gamma_{\bar x, \bar \theta}).
\end{align}
It is clear that by construction $\partial_{\bar\theta} \gamma_{\bar x, \bar \theta}; \partial_t \gamma_{\bar x, \bar \theta}$ just provides the measure of the volume form of the disc generated by the normal geodesics from $p$, say $dVol_{D_{\bar x}}(\bar \theta,t)$.
Notice that for generic $M$ it is expected to depend on $\bar x$. The remaining contribution can be computed by employing the Jacobi equation w.r.t. the Fermi frame. It is (the tilde just means we are restricting to the directions of the Fermi frame)
\begin{align}
 \ddot {\tilde J}_{ab} +\sum_c R(\dot \gamma, e_a, \dot \gamma, e_c) \tilde J_{cb}=0, \label{Jacobian-equation}
\end{align}
where the dot indicates derivative w.r.t. to $t$. Here $\gamma\equiv \gamma_{\bar x, \bar \theta}$ and $e_a$ are to be intended as Fermi transported along the geodesic. Because of this, we have
\begin{align}
 \frac d{dt} R(\dot \gamma, e_a, \dot \gamma, e_c)=(\nabla_{\dot \gamma} R)(\dot \gamma, e_a, \dot \gamma, e_c).
\end{align}
Notice that for $N$ compact the matrix $R(\dot \gamma, e_a, \dot \gamma, e_c)$ is symmetric and positive definite. In general, the solution of equation \eqref{Jacobian-equation} is completely determined by the Cauchy data $\tilde J_{cb}|_{t=0}$,
$\dot {\tilde J}_{cb}|_{t=0}$. For example, deriving \eqref{Jacobian-equation} in $t=0$, we get
\begin{align}
 \dddot {\tilde J}_{ab}+\sum_c R(\dot \gamma, e_a, \dot \gamma, e_c) \dot{\tilde J}_{cb} +\sum_c (\nabla_{\dot \gamma} R)(\dot \gamma, e_a, \dot \gamma, e_c) \tilde J_{cb}=0,
\end{align}
and iterating this operation, one gets $\left. \frac {d^n {\tilde J}_{ab}}{dt^n} \right|_{t=0}$ as a function of $\tilde J_{ab}$, $\dot {\tilde J}_{ab}$ and $R(\dot \gamma, e_a, \dot \gamma, e_c)$ and all its covariant derivatives up to order $n-2$ along the direction $\dot \gamma$, in the point $p$. More in general, we can write formally the solution in the form
\begin{align}
 \tilde J(\bar x, \bar \theta, t)=J_0(\bar x, \bar \theta)+\sum_{j=1}^\infty (A_j  J_0 +B_j \dot J_0) (\bar x, \bar \theta)  \frac {t^j}{j!}, \label{general-jacobian}
\end{align}
where $J_0$, $\dot J_0$, $A_j$, $B_j $ are $n\times n$ matrix-valued functions of $(\bar x, \bar \theta, t)$ defined by
\begin{align}
 J_0=&\tilde J(\bar x, \bar \theta, 0), \\
 \dot J_0=& \frac {d\tilde J}{dt} (\bar x, \bar \theta, 0), \\
 (A_1)_{ab}=&0, \qquad (B_1)_{ab}=\delta_{ab}, \\
 A_{j+1}=& \nabla_{\dot \gamma} (A_j)_{ab}-\sum_c (B_j)_{ac}R(\dot \gamma, e_c, \dot \gamma, e_b), \\
 B_{j+1}=& \nabla_{\dot \gamma} (B_j)_{ab}+ (A_j)_{ab}.
\end{align}
Let us assume the analyticity condition that for any given $(\bar x, \bar \theta)$ the series \eqref{general-jacobian} has a strictly positive convergence radius. Since $N$ is compact and $M$ is closed, then also $M$ is compact and there is a minimum positive
radius $\tau$, such that \eqref{general-jacobian} converges uniformly in any region $t\leq \varepsilon <\tau$. We can fix such an $\varepsilon$ to define the tube.
Moreover, notice that $\tilde J(\bar x, \bar \theta, 0)$ determines the change of variables along $M$, and its determinant does not depend on $\bar \theta$, while
\begin{align}
 \dot {\tilde J} (\bar x, \bar \theta, 0)_{ab}=\sum_{c=1}^n \sum_{s=1}^q \omega^s(\bar\theta) K^s_{ac}(\bar x) {\tilde J} (\bar x, \bar \theta, 0)_{cb},
\end{align}
where $K^j_{ab}$ is the second fundamental form of the embedding of $M$ along the direction $n^j$ and $\omega^j$ are the director cosines defined above.
Therefore, we have proven the following proposition:
\begin{prop}
Let $M$ be a Riemannian submanifold of codimension $q$ in a compact Riemannian manifold $N$ of dimension $q+n$. Let $M_\varepsilon$ be a tube of radius $\varepsilon$ around $M$, coordinatized as above. Then, the volume element $dV$ in the tube is
\begin{align}
 dV= dVol_{D_{\bar x}}(\bar \theta,t) dVol_M(\bar x) \det\left( I_n+\sum_{j=1}^\infty \left(A_j (\bar x, \bar \theta)  +B_j(\bar x, \bar \theta)  \sum_{s=1}^q \omega^s(\bar\theta) K^s(\bar x)\right)  \frac {t^j}{j!},  \right),\label{formulone}
\end{align}
where $K^s$ is the symmetric matrix with components $K^s_{ab}$, that is the second fundamental form along the normal direction $s$, $I_n$ is the $n\times n$ identity matrix.
\end{prop}
This very general formula is clearly of poor practical usage since for applications one needs to have control of the Riemann tensor and all its covariant derivatives. In any case, we can see that if we look for the concentration of the measure around $M$, the main ingredients entering into the game are the extrinsic curvatures $K^s(\bar x)$ and the volume of $M$. Of course, large values
of the curvatures may amplify the last factor. However, if the sign of the curvatures is constant because we are looking for convex regions, then, large values of the curvatures may correspond to a small value of the volume of $M$. For example, it is well
known that the measure of the spheres concentrates on equators which are totally geodesic subvarieties, thus having zero extrinsic curvatures. This maximizes the volume of $M$.  \\
This suggests that the best candidates for the concentration of the measure are totally geodesic subvarieties. The contribution of the curvatures should then be to maximize the dependence on $t$ in $t=0$ through the coefficients $A_j$. However, it is
quite hard to say more in the general case, both because it is not guaranteed the existence of totally geodesic subvarieties and because it is quite hard to have a uniform control on the coefficients $A_j$ and $B_j$. For these reasons, we now move to specific examples.

\subsection{Compact Symmetric spaces}
We want to apply our general formula to the case of compact symmetric spaces $\Sigma= G/H$, where $G$ is a compact Lie group and $H$ is a symmetrically embedded subgroup. The reason is that they are simple enough to allow for a very explicit calculation of the coefficients $A_j$ and $B_j$, and, at the same time, they contain several totally geodesic submanifolds.  In a sense, they are the simplest generalizations of $S^N =SO(N+1)/SO(N)$. Here we consider the case where $G$. is a
simple group, but our construction can be extended to semisimple groups in an obvious way. We assume that the dimension of $\Sigma$ is $N=n+q$ while $M$ is an $n$ dimensional submanifold.
$\Sigma$ is endowed by a metric $g_{ij}$ that is invariant under both the left and the right translations generated by $G$. Since $G$ is compact, the metric is induced by the Killing form of $Lie(G)$, up to a (negative) constant. The corresponding
Riemann tensor is covariantly constant. In particular, $\Sigma$ is an Einstein manifold with Ricci tensor $R_{ij}=\frac SN g_{ij}$, where the scalar curvature $S$ is a constant. We have the following orthogonal decomposition
\begin{align}
 Lie (G)=Lie(H)\oplus Lie(H)^\perp.
\end{align}
The elements of $Lie(H)$ act as infinitesimal isometries leaving fixed the points of $\Sigma$, while the elements of $Lie(H)^\perp$ generate translations. At each point $p$ of $M$ we can take of vectors $\{\vec n_1,\ldots, \vec n_q\}$ forming a basis
of the normal space of $T_pM$ in $T_p\Sigma$. We can take $\vec n_j$ as elements of $Lie(H)^\perp$. The disc of radius $a$ in $T_p M^\perp$ defined by
\begin{align}
 D_a(p)=\{ x_1 \vec n_1+\cdots+x_q \vec n_q| x_1^2+\cdots+x_q^2\leq a^2 \}
\end{align}
is mapped to a geodesic disc by the exponential map. The orthogonal geodesics from $p$ are thus of the form
\begin{align}
\gamma(t)= e^{t \sum_j v^j \vec n_j}\cdot p
\end{align}
where $\cdot$ indicates the action of the elements of $G$ on $\Sigma$, $\sum_j (v^j)^2=1$, and the exponential is in the sense of groups. Then, $\varepsilon$ is the geodesic distance of $\gamma(a)$ from $p$.\\
The main point now is that, since the Riemann tensor is covariantly constant, we have
\begin{align}
 (\nabla_{\dot \gamma} R)(\dot \gamma, e_a, \dot \gamma, e_c)=0.
\end{align}
Therefore, the matrix
$$R(\dot \gamma, e_a, \dot \gamma, e_c)=R(n(\bar x;\bar\theta), e_a, n(\bar x;\bar\theta), e_c)_{\bar x}=: A(\bar x, \bar \theta)_{ac}$$
is constant in $t$ and we need just to evaluate it in the point $p$. Moreover, $\Sigma$ is compact so that the matrix $A$ is symmetric and positive definite. Hence, it exists an orthogonal matrix $\Omega(\bar x, \bar \theta)$ such that
\begin{align}
 A(\bar x, \bar \theta)=\Omega(\bar x, \bar \theta) D^2(\bar x, \bar \theta) \Omega(\bar x, \bar \theta)^T,
\end{align}
where $D^2$ is a diagonal matrix with positive eigenvalues $d_j^2(\bar x, \bar \theta)$.
If, for any given real function $f$, we define $f(Dt)$ as the diagonal matrix having $f(d_j t)$ as diagonal elements, and $f(\sqrt A t)=\Omega f(Dt) \Omega^T$, then we get for the volume element $dV$ in the tube is given by the following
\begin{prop}
Let $M$ be a Riemannian submanifold of codimension $q$ in a compact Riemannian symmetric manifold $\Sigma$ of dimension $q+n$. Let $M_\varepsilon$ be a tube of radius $\varepsilon$ around $M$, coordinatized as above. Then, the volume element $dV$ in the tube is
\begin{align}
 dV= dVol_{D_{\bar x}}(\bar \theta,t) dVol_M(\bar x) \det\left( \cos (\sqrt {A(\bar x, \bar \theta)} t) + \frac {\sin (\sqrt {A(\bar x, \bar \theta)} t)}{\sqrt {A(\bar x, \bar \theta)}}  \sum_{s=1}^q \omega^s(\bar\theta) K^s(\bar x)\right). \label{formulozza}
\end{align}
\end{prop}
This is what we get by a direct application of \eqref{formulone} with a constant matrix $R$. Notice that in general, the volume element of the disc depends on its center $\bar x$, as well as the matrix $A(\bar x, \bar \theta)$.\\
 We now restrict further ourselves to the case when $M$ is a totally geodesic submanifold of $\Sigma$. In this case $K^s(\bar x)=0$.
 Since the sub-manifold $\Sigma$ is totally geodetic, any two points in $M$ are connected by a geodetic of $\Sigma$ which is also a geodesic for $M$. Moreover, $\Sigma$ is symmetric, so it has covariantly constant Riemann tensor (and metric, obviously). Then $dVol_{D_{\bar x}}(\bar \theta,t)$ and $A(\bar x, \bar \theta)$ are independent on $\bar x$ and the formula further reduces to
 \begin{align}
 dV= dVol_{D}(\bar \theta,t) dVol_M(\bar x) \prod_{a=1}^n \cos (d_a(\bar \theta) t). \label{formulino}
\end{align}
Notice that for a compact symmetric space, analyticity is guaranteed, and $t$ can thus be extended so that the above parametrization covers the whole manifold up to a measure null set (the set of focal points). Therefore, we can prove the following proposition.
We can now notice that the range is such that the cosine factors remain non-negative. In this case, we can notice that for $x\in [0,\pi]$ one has
\begin{align}
 \cos x\leq e^{-\frac {x^2}2},
\end{align}
from which we get that everywhere
\begin{align}
 0\leq \prod_{a=1}^n \cos (d_a(\bar \theta) t) \leq e^{-\frac {t^2}2\sum_{a=1}^n d^2_a(\bar \theta)} \label{138}
\end{align}
\begin{prop}
Let $(M_n,\Sigma_n)$ be a family, labeled by $n$, of $n$-dimensional totally geodesic submanifolds $M_n$ of symmetric spaces $\Sigma_n$ of dimension $n+1$ and constant diameter. Let $M^{\varepsilon_n}_n$ be the tube of geodesic radius $\varepsilon_n$ centered in $M_n$.
 Finally, let $m_n$ the Riemannian measure over $\Sigma_n$ normalized so that $\mu_n(\Sigma_n)=1$. If $\lim_{n\to\infty} \sqrt n\varepsilon_n=\infty$ then
 \begin{align}
 \lim_{n\to\infty} \mu(\Sigma-M^{\varepsilon_n}_n)=0.
 \end{align}
\end{prop}
\begin{proof}
 Suppose we consider the measure \eqref{formulino} normalised to 1. We also assume that $\Sigma_n$ has diameter $L$. We need to consider
 \begin{align}
 \lim_{n\to\infty} \int_{\Sigma_n}  dVol_{D}(\bar \theta,t) dVol_M(\bar x) f_n
 \end{align}
 where
 \begin{align}
 f_n=\prod_{a=1}^n \cos (d_a(\bar \theta) t) \chi_{(\Sigma-M^{\varepsilon_n}_n)},
 \end{align}
and $\chi_E$ is the characteristic function of the set $E$. Since the codimension of $\Sigma$ is 1, we have that
\begin{align*}
\sum_{a=1}^n d^2_a(\bar \theta)= \sum_{a=1}^n R(n(\bar x;\bar\theta), e_a, n(\bar x;\bar\theta), e_c)_{\bar x}=Ric(n(\bar x;\bar\theta), n(\bar x;\bar\theta)),
\end{align*}
where $Ric$ is the Ricci tensor of $M$. It is known that for a symmetric manifold of constant diameter and dimension $s$, the Ricci tensor has the form $R=(as+b)g$, $g$ being the invariant metric and $a>0$ and $b$ constants independent on $s$. For example, this follows easily from the calculations in \cite{CU22} and \cite{CaSc22}. Applied to our case and using \eqref{138} this shows that
\begin{align}
 f_n\leq e^{-\frac {t^2}2 (an+b)} \chi_{(\Sigma-M^{\varepsilon_n}_n)}
\end{align}
for some constants $a>0$ and $b$. Since the integral is extended in $t\geq \varepsilon_n$, we get that the integrand goes to zero uniformly at worst as $e^{-\frac a2 \varepsilon^2_n n}$ when $n$ diverges. Since $\lim_{n\to\infty} \varepsilon^2_n n=0$,
the assert is proved.
\end{proof}
{\bf Remark:} The assumption for the codimension to be 1 has been made to keep the proof technically simple. We believe that the same proposition is true for constant codimension $q>1$ and we expect it to hold also for codimension $q_n$, if it doesn't grow
too fast with $n$. We leave the investigation of this point for future work.

\section{Characterization through wasserstein and box distance}\label{box}

 Let $(M,\mu,g)$ be a Riemann manifold and $C\subseteq M$ a submanifold. The projection map $proj_C:M\rightarrow C$ is defined as a map such that $d(x,proj_C(x))=d(x,C)$, where $d$ is the geodetic distance associated with $g$. For the existence and, more in general, a theory of distance functions in $\mathbb{R}^n$ context see e.g. \cite{CanSin}; for an extension of this theory to a Riemannian context see e.g.\cite{ManteMen}, \cite{Fath}.

\begin{prop}\label{WassConcLoc}
 Let $((N_n,\mu_n,g_n),M_n)$ a sequence of Riemannian manifolds, with $M_n \subsetneq N_n$ Riemannian submanifold with endowed measure $\#proj_{N_n,M_n}\mu_n$. Let $d_{W_2}$ be the Wasserstein distance of order 2.
 If $d_{W_2}(\mu_n,\#proj_{N_n,M_n}\mu_n)\rightarrow 0$ then $M_n$ is a concentration locus for $N_n$
\end{prop}
\begin{proof}
If $\pi_n$ is the geodesic projection on $M_n$ and $d_n$ the geodesic distance generated by the invariant metric $g_n$, then, the cost to transport the mass $m_n=\mu_n(N_n\setminus M_n^{\varepsilon_n})$ in terms of $d_{W_2}^2$ is at least
\begin{align*}
\int_{N_n\setminus M_n^{\varepsilon_n}} d_n^2(x,\pi_n(x)) d\mu_n(x)>\varepsilon_n^2 m_n,
\end{align*}
for $M_n^{\varepsilon_n}$ a tubular neighbourhood of radius $\varepsilon_n$ of $M_n$. In particular, for any fixed choice $\varepsilon_n=\varepsilon>0$, since $d_{W_2}(\mu_n,\#proj_{N_n,M_n}\mu_n)\rightarrow 0$, we get that $m_n\rightarrow 0$.
\end{proof}

Proposition \ref{WassConcLoc} cannot be reversed.
Indeed, in case of a concentration locus $M_n\subseteq N_n$, if the diameter of $N_n$ is unbounded $d_W$ could not converge to 0.

Nevertheless the following holds.
\begin{prop}\label{ConcW1}
  Let $(N_n,M_n)$ a sequence as in the previous Proposition such that volumes and diameters of $N_n$ are bounded by $h>0$. If $M_n$ is a concentration locus for $N_n$ then for all $n\in\mathbb{N}$ $d_{W_1}(\mu_n,\#proj\mu_n)\rightarrow 0$
\end{prop}

\begin{proof}
  By contradiction, suppose that for infinitely many $n$ $d_{W_1}(\mu_n,\#proj\mu_n)\ge k>0$. Since the distance function determines the optimal transport through the project function, $\int_{N_n}d_n(x,\pi_n(x)) d\mu_n(x) \ge k$, where $\pi_n$ is the geodetic
  projection over $M_n$ and $d_n$ the geodetic distance generated by the invariant metric $g_n$ on $N_n$.
Since $M_n$ is a concentration locus for $N_n$ we can set $n'$ in such a way for all $n>n'$ $\epsilon_n, \varepsilon_n<\frac{k}{4h}$ and, $\mu_n(N_n\setminus M_n^{\varepsilon_n})<\epsilon_n$. Therefore,
\begin{align*}
 k&\le\int_{N_n\setminus M_n^{\epsilon_n}}d_n(x,\pi_n(x))d\mu_n(x) +\int_{M_n^{\epsilon_n}}d_n(x,\pi_n(x))d\mu_n(x)\cr
 &\le\int_{N_n\setminus M_n^{\epsilon_n}}hd\mu_n(x)+\int_{M_n^{\epsilon_n}}\frac{k}{4h}d\mu_n(x)\le 2h\frac{k}{4h}=\frac k2,
\end{align*}
 which is a contradiction.

\end{proof}

\begin{rem}\label{distance}
Observe that convergence in Wasserstein distance (Strassen's Theorem \cite{Shi} implies Prohorov distance convergence) implies box distance convergence (Proposition 4.12 \cite{Shi}), which in turn implies $d_{conc}$ convergence (Proposition 5.5 (2) \cite{Shi}).

Therefore, let $M$ be a complete separable metric space
and ($\mu_i$) a sequence of Borel probability measures on $M$.
Consider the following three conditions:
(1) $\mu_i$ converges weakly to $\mu$.
(2) $(M,\mu_i)$ box-converges to $(M,\mu)$.
(3) $(M,\mu_i)$ $d_{conc}$-converges (or concentrates) to $(M,\mu)$.

Then, the following implications hold:
(1) $\Rightarrow$ (2) $\Rightarrow$ (3).

A counterexample of (2) $\Rightarrow$ (1) is easy.
Just take a sequence $x_i$ in $M$ and let $\mu_i$ be the Dirac measure at $x_i$.
Then, all $(M,\mu_i)$ are mm-isomorphic to each other, so that (2) holds.
However (1) does not hold if $x_i$ does not converge in $M$.

A counterexample of (3) $\Rightarrow$ (2) is the sequence of unit spheres $S^n(1)$
with dimension $n$ going to infinity.  $S^n(1)$ $d_{conc}$-converges to one-point space,
but it is divergent for box-distance (see Cor. 5.20 \cite{Shi}).

\end{rem}

\begin{teo}\label{ConcLocDconc}
  Let $(N_n,\mu_n,g_n)$ a sequence of Riemann manifolds with haar measure $\mu_n$, geodesic distance $d_n$ generated by the invariant metric $g_n$ (they are in particular mm-spaces), and $M_n\subseteq N_n$ sub manifolds with measures $\sigma_n=\#proj\mu_n$ and metrics $g'_n=g_n|_{M_n}$. \\
Provided that $d_{box}((N_n,\mu_n),(M_n,\sigma_n))\rightarrow 0$ and assuming $(M_n,\sigma_n,g'_n) \rightarrow_ {d_{conc}} (M,\sigma,g')$ then\\ $(N_n,\mu_n,g_n) \rightarrow_ {d_{conc}} (M,\sigma,g')$.
\end{teo}
\begin{proof}
 By Proposition 5.5 \cite{Shi}, $d_{conc}((N_n,\mu_n,g_n),(M_n,\sigma_n,g'_n))\le d_{box}((N_n,\mu_n,g_n),(M_n,\sigma_n,g'_n))$. Hence, $d_{conc}((N_n,\mu_n,g_n),(M_n,\sigma_n,g'_n))\rightarrow 0$.\\
 The following chain of inequalities
 $$d_{conc}((N_n,\mu_n,g_n),(M,\sigma,g))\le d_{conc}((N_n,\mu_n,g_n),(M_n,\sigma_n,g'_n))\ +\ d_{conc}((M_n,\mu_n,g_n),(M,\sigma,g))$$
 plainly drives to the thesis.
\end{proof}

\begin{rem}
The above Theorem holds also with the hypothesis $d_{W_1}(\mu_n,\sigma_n)\rightarrow 0$.
Indeed, we can consider $\sigma_n$ as a measure $\sigma '_n$ on $N_n$ having support $spt(\sigma'_n)=M_n$ where it is equal to $\sigma_n$. Without loss in generality, we can consider $(N_n,\sigma'_n,g_n)$ equal, as mm-spaces, to $(M_n,\sigma_n,g'_n)$, since they are mm-isomorphic. Indeed, two mm-spaces are mm-isomorphic if there exists an isometry between their supports of their respective measures (see \cite{Gro99} p.117).

\end{rem}

Now the following Corollary follows

\begin{cor}
 Let $(N_n,M_n)$ a sequence of Riemannian manifolds $M_n\subsetneq N_n$ such that volumes and diameters of $N_n$ are bounded by $h\in\mathbb{R}$. If $M_n$ is a concentration locus for $N_n$ and $ (M_n,\sigma_n,g'_n)\rightarrow_ {d_{conc}} (M,\sigma,g')$ then $(N_n,\mu_n,g_n) \rightarrow_ {d_{conc}} (M,\sigma,g')$.
\end{cor}

Now we move towards a $d_{box}$-characterization.

\begin{prop}\label{concbox}
Let $M_n$ be a concentration locus for $N_n$, $M_n$ totally geodesic submanifolds, with the properties that the geodesic distance on $M_n$ is the same as in $N_n$. 
 Then, $d_{box}(N_n,M_n)$ converges to 0.
\end{prop}

\begin{proof}
   Let us first observe that if $M_n^{\epsilon_n}$ is the tubular neighbourhood of radius $\epsilon_n$, then, for $s,t\in M_n^{\epsilon_n}$ we have
     $$\vert d_n(s,t)- d'_n(proj_{M_n}(s),proj_{M_n}(t))\vert \le O(\epsilon_n),$$
     where $d_n'$ is the geodesic distance in $M_n$.
  Indeed, inside the tube by elementary distance inequalities $\vert d_n(s,t)- d_n(proj_{M_n}(s),proj_{M_n}(t))\vert \le 2\epsilon$, and since by hypothesis $d_n(proj_{M_n}(s),proj_{M_n}(t))=d'_n(proj_{M_n}(s),proj_{M_n}(t))$.\\
   Now define two parameters, $\phi,\psi$ for $N_n$ and $M_n$, respectively,
   $\phi:[0,\epsilon_n)\rightarrow N_n\setminus M_n^{\epsilon_n}$ and $\phi:[\epsilon_n,1)\rightarrow  M_n^{\epsilon_n}$, where $\epsilon_n$ is chosen so that $\mu_n(N_n\setminus M_n^{\epsilon_n})\le\epsilon_n$. This is always possible, since if we have 
   two sequences $a_n$ and $b_n$ positive, converging to zero and such that $\mu_n(N_n\setminus M_n^{a_n})\le b_n$, then we can choose $\epsilon_n=\max \{a_n,b_n\}$.
   Let $\psi:=proj_{M_n}\circ\phi$.\\ 
   From the above inequality, it follows:
   \begin{align*}
   \vert d_n(\phi(s),\phi(t))-  d_n'(\psi(s),\psi(t))\vert\le O(\epsilon_n).
   \end{align*}
   This implies the thesis.

\end{proof}

Combining Proposition \ref{concbox} and Theorem \ref{ConcLocDconc} we get the following Corollary

\begin{cor}\label{ConcLocConv}
 Let $(N_n,\mu_n,g_n)$ be a sequence of mm-spaces, with $M_n$ totally geodesic and with the same geodesic distance as in $N_n$. Suppose that $M_n$ define a concentration locus for $N_n$ with a sequence of radii $\epsilon_n\rightarrow 0$, and $(M_n,\sigma_n,g'_n) \rightarrow_ {d_{conc}} (M,\sigma,g')$. Then, $(N_n,\mu_n,g_n) \rightarrow_ {d_{conc}} (M,\sigma,g')$.
\end{cor}

In \cite{CU22}, by using the Macdonald formula \cite{Ma}, it is shown that $SU(n), Spin(n), Usp(n)$ have all totally geodesic concentration loci.

Observe that even under hypothesis $d_{box} (N_n,H_n)\rightarrow 0$, for any $N_n \in SO(n),SU(n),Spin(n)$, by Corollary 5.20 \cite{Shi}, all of them don't determine $d_{box}$ divergent sequences.

Observe that Corollary \ref{ConcLocConv} applies even if $N_n$ is not a Levy family. In this case, the sequence concentrates anyway to $M$ which is not necessarily a point.
This unveils that the concentration phenomenon is far from being exhausted by the Levy families. 
In particular, if $N_n$ are topological groups, by a Schneider result \cite{Sc}, $M$ should be not only the concentration set but also an $N$-invariant subspace of $S(N)$, where $N$ is a second-countable topological group completion of $\bigcup N_n$, and 
$S(N)$ is the Samuel compactification of $N$.
If $M$ is minimal, then it is the universal minimal flow of $N$.
Since concretely describable universal minimal flows are rather rare, this could be a way to construct them.\\
These constructions can find interesting applications to the sequences of $U(N)$ with different rescaled geometries.

\section*{Acknowledgments}
We are extremely grateful to Alessio Figalli for his insightful suggestions and precise remarks. We also thank Carlo Mantegazza and Takeshi Shioya for some helpful discussions.

\end{document}